Compute the mean and variance of the following discrete probability distribution: (Round the final answers to 2 decimal places.)

X P(X)
4 0.4
6 0.1
10 0.5


Mean =
Variance =
1
Expert's answer
2020-03-26T14:40:09-0400

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 X & P(X) \\ \hline
 4 & 0.4 \\
 \hdashline
 6 & 0.1 \\
 \hdashline
 10 & 0.5 
\end{array}"

"\mu =\sum\limits ^3_{i=1}p(x_i)x_i=0.4\times 4+0.1\times 6+0.5\times 10=7.2"

"\sigma ^2= \sum\limits ^3_{i=1}p(x_i)(x_i-\mu)^2=0.4\times (-3.2)^2+0.1\times (-1.2)^2+0.5\times 2.8^2=8.16"


Answer: mean = 7.2, variance = 8.16.